$SPAD/src/input farray.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool farray.output
)set message test on
)set message auto off
)clear all
--S 1 of 16
flexibleArray [i for i in 1..6]
--R
--R
--R   (1)  [1,2,3,4,5,6]
--R                                         Type: FlexibleArray(PositiveInteger)
--E 1
--S 2 of 16
f : FARRAY INT := new(6,0)
--R
--R
--R   (2)  [0,0,0,0,0,0]
--R                                                 Type: FlexibleArray(Integer)
--E 2
--S 3 of 16
for i in 1..6 repeat f.i := i; f
--R
--R
--R   (3)  [1,2,3,4,5,6]
--R                                                 Type: FlexibleArray(Integer)
--E 3
--S 4 of 16
physicalLength f
--R
--R
--R   (4)  6
--R                                                        Type: PositiveInteger
--E 4
--S 5 of 16
concat!(f,11)

--R
--R
--R   (5)  [1,2,3,4,5,6,11]
--R                                                 Type: FlexibleArray(Integer)
--E 5
--S 6 of 16
physicalLength f
--R
--R
--R   (6)  10
--R                                                        Type: PositiveInteger
--E 6
--S 7 of 16
physicalLength!(f,15)
--R
--R
--R   (7)  [1,2,3,4,5,6,11]
--R                                                 Type: FlexibleArray(Integer)
--E 7
--S 8 of 16
concat!(f,f)
--R
--R
--R   (8)  [1,2,3,4,5,6,11,1,2,3,4,5,6,11]
--R                                                 Type: FlexibleArray(Integer)
--E 8
--S 9 of 16
insert!(22,f,1)
--R
--R
--R   (9)  [22,1,2,3,4,5,6,11,1,2,3,4,5,6,11]
--R                                                 Type: FlexibleArray(Integer)
--E 9
--S 10 of 16
g := f(10..)
--R
--R
--R   (10)  [2,3,4,5,6,11]
--R                                                 Type: FlexibleArray(Integer)
--E 10
--S 11 of 16
insert!(g,f,1)
--R
--R

--R   (11)  [2,3,4,5,6,11,22,1,2,3,4,5,6,11,1,2,3,4,5,6,11]
--R                                                 Type: FlexibleArray(Integer)
--E 11
--S 12 of 16
merge!(sort! f, sort! g)
--R
--R
--R   (12)  [1,1,2,2,2,2,3,3,3,3,4,4,4,4,5,5,5,5,6,6,6,6,11,11,11,11,22]
--R                                                 Type: FlexibleArray(Integer)
--E 12
--S 13 of 16
removeDuplicates! f
--R
--R
--R   (13)  [1,2,3,4,5,6,11,22]
--R                                                 Type: FlexibleArray(Integer)
--E 13
--S 14 of 16
select!(i +-> even? i,f)
--R
--R
--R   (14)  [2,4,6,22]
--R                                                 Type: FlexibleArray(Integer)
--E 14
--S 15 of 16
physicalLength f
--R
--R
--R   (15)  8
--R                                                        Type: PositiveInteger
--E 15
--S 16 of 16
shrinkable(false)$FlexibleArray(Integer)
--R
--R
--R   (16)  true
--R                                                                Type: Boolean
--E 16
)spool
)lisp (bye)

———-

References^{[1] nothing}

6